OpenStudy (indiiavery):

If x varies inversely with y and x = 8 when y = 6, find y when x = 10. Answers A:y=5/12 B:y=7.5 C:y=4.8 D:y=40/3 Please help. Also I need help figuring out how to solve it. (Like working it out etc.)

OpenStudy (anonymous):

Do you know what x varying inversely with y means?

OpenStudy (indiiavery):

No, having trouble understanding it. :/

OpenStudy (anonymous):

well it means that x / y = A. Where A is a constant. So you have x / y = x / y So 8 / 6 = 10 / y

OpenStudy (anonymous):

Do you understand now?

OpenStudy (indiiavery):

Oh okay, so would I have to like Multiply or Divide?

OpenStudy (indiiavery):

Yes.

OpenStudy (anonymous):

You first need to get the y on top, you can do that by 1) multiplying by y on both sides and then dividing by 8 and multiplying by 6, or 2) you can take the inverse of both sides and multiply by 10.

OpenStudy (indiiavery):

Ohhh okay, thank you so much for taking your time to explain and helping me with my problem. :)

OpenStudy (anonymous):

No problem :)